An oil tanker can be emptied by the main pump in 5hours. An auxillary pump can empty the tanker in 9 hours. If the main pump is started at 5PM, when should the auxillary pump be started so the tanker is emptied by 9PM?
How do I determine the "time"? I get to 1/9t=1/5 and not sure what to do.
Thank you.

Answer by nerdybill(7384)

An oil tanker can be emptied by the main pump in 5hours. An auxillary pump can empty the tanker in 9 hours. If the main pump is started at 5PM, when should the auxillary pump be started so the tanker is emptied by 9PM?
.
From 5pm to 9pm is 4 hours...
.
Let x = number of hours to run aux pump
then
4(1/5) + x(1/9) = 1
4/5 + x/9 = 1
multiplying both sides by 45:
36 + 5x = 45
5x = 9
x = 9/5 hours
.
9/5 hour is the same as
9/5 * 60 = 108 minutes
1 hour and 48 minutes (time needed to run aux pump)
.
Time to start aux pump:
9 pm minus 1 hour and 48 minutes:
7:12 PM